# Secure IVSHMEM: End-to-End Shared-Memory Protocol with Hypervisor-CA Handshake and In-Kernel Access Control

### v1

@@
Abstract
In-host shared memory (IVSHMEM) enables high-throughput, zero-copy communication between virtual machines, but today’s implementations lack any security control, allowing any application to eavesdrop or tamper with the IVSHMEM region. This paper presents Secure IVSHMEM, a protocol that provides end-to-end mutual authentication and fine-grained access enforcement with negligible performance cost. We combine three techniques to ensure security: (1) channel separation and kernel module access control, (2) hypervisor-mediated handshake for end-to-end service authentication, and (3) application-level integration for abstraction and performance mitigation. In microbenchmarks, Secure IVSHMEM completes its one-time handshake in under 200 $\mu$s and sustains data-plane round-trip latencies within 5% of the unmodified baseline, with negligible bandwidth overhead. We believe this design is ideally suited for safety- and latency-critical in-host domains, such as automotive systems, where both performance and security are paramount.
Index Terms:
IVSHMEM, Inter-VM Shared Memory, End-to-End Security
I Introduction
@@
The automotive industry is rapidly evolving, driven by advances in semiconductor technology that have shifted system architectures from traditional microcontrollers (MCUs) to powerful Systems-on-Chip (SoCs). This evolution not only enhances computational capabilities but also paves the way for Software-Defined Vehicles (SDVs), where flexibility, scalability, and rapid updates are paramount. In SDVs, virtualization technology plays a crucial role by enabling the coexistence of multiple virtual machines (VMs) on a single hardware platform, ensuring isolated yet efficient execution of diverse applications. For example, modern cockpit domain controllers often deploy separate VMs for real-time operations (RTOS) and infotainment systems, which is essential for balancing performance and safety[1][2].
@@
Inter-VM communication in these environments is critical. Traditional approaches, such as TCP/UDP over a network stack or even UART-based messaging, often fall short in terms of speed and resource efficiency[3]. Alternative solutions like VirtIO offer a para-virtualized communication mechanism through ring buffers (VirtQueues), but they do not fully leverage the benefits of shared physical memory.[4] IVSHMEM (Inter-VM Shared Memory) addresses these limitations by mapping each VM’s virtualized PCI device to a common physical memory region, allowing rapid data exchange through shared memory[5][6][3]. Despite its performance advantages, this method introduces significant security challenges; multiple VMs accessing the same memory space creates vulnerabilities where a compromised or malicious VM could potentially access or modify data belonging to another VM[7].
@@
This concern is particularly acute in scenarios where critical systems interact with less secure environments. For instance, when an RTOS communicates with an Android-based infotainment VM, there is a tangible risk that a malicious application within Android might tamper with the shared memory region[2]. Such tampering could result in attacks ranging from man-in-the-middle to eavesdropping, ultimately compromising system stability and safety.
@@
II Background
In this section, we examine the IVSHMEM (Inter-VM Shared Memory) mechanism, outline its challenges, and review recent progress.
II-A IVSHMEM: Mechanism and Architecture
IVSHMEM is a specialized implementation of shared memory IPC designed for virtualized environments. It emulates a virtual PCI device to expose the shared memory’s base address and size to guest VMs[8][9]. As shown in Figure 1, the hypervisor exposes the shared-memory region to the guest VM, and the UIO kernel driver maps the emulated PCI device, allowing applications to access that memory directly via /dev/uioX. The design leverages the standardized PCI configuration to facilitate memory mapping and efficient communication. Specifically, IVSHMEM utilizes:
@@ II-A IVSHMEM: Mechanism and Architecture
The doorbell interrupt mechanism enabled by this configuration allows VMs to notify one another when new data is available, ensuring efficient core utilization and reducing latency in inter-VM communication[10].
Figure 1: The overview of IVSHMEM architecture
II-B Security Concerns of Shared Memory
While shared memory IPC within a single operating system benefits from well-established security mechanisms—such as file access controls, sandboxing, and enhanced security modules like SELinux or AppArmor—IVSHMEM presents unique challenges[11][12]. In a traditional OS environment, the operating system enforces strict access controls over shared memory regions, ensuring that only authorized processes can read or write data. However, when multiple, potentially untrusted VMs share the same memory space, these protections are significantly diminished. A compromised or malicious VM could easily access or tamper with data in the shared region, leading to unauthorized data disclosure, corruption, or even system instability[7].
II-C Secure Communication Over Insecure Channels
The challenge of ensuring secure communication in IVSHMEM environments is analogous to securing communication over the Internet, where multiple parties exchange information over an inherently insecure channel. In network communications, protocols such as TLS rely on key exchange mechanisms, mutual authentication, and end-to-end encryption to safeguard data integrity and confidentiality[13][14]. Similarly, secure multi-party communication techniques—such as Diffie-Hellman key exchange and advanced encryption standards—are employed to establish trust even when the channel is compromised[15][16].
In the context of IVSHMEM, the situation is even more complex because multiple services must share the same restricted memory space as a communication channel. This necessitates the design of a secure protocol that not only ensures confidentiality and integrity—akin to TLS or other network security protocols—but also accommodates the shared nature of the memory resource. Our research addresses these challenges by proposing a secure protocol that protect the data transmitted via IVSHMEM, while also mitigating the performance overhead typically associated with such security measures.
II-D Recent Progress of IVSHMEM Communication
@@
The SIVSHM project introduces a segmentation approach to IVSHMEM, enhancing security by isolating shared memory regions among VMs. However, this strict isolation introduces notable overheads from reduced buffer size. [7].
Performance-centric solutions, such as XenLoop and MemPipe, have been proposed to optimize IVSHMEM by improving transparency and reducing latency[17][3][18][19][20][21]. These solutions integrate seamlessly with traditional socket-based network stacks, allowing applications to benefit from high-speed shared memory communication without explicit changes. However, these solutions primarily prioritize performance and transparency and lack of mechanisms for security threats[3].
Other approaches have leveraged hypervisor-managed policies and features, like Xen’s grant tables, to enforce finer-grained security controls. Grant tables establish explicit, controlled memory-sharing agreements between VMs, restricting access to designated regions. However, this technique is tied to a specific hypervisor and does not provide a general-purpose driver interface [22].
@@ III-D Security Goals
III-E Limitations of Conventional Security Protocols
@@
  1. 1.
    Performance Overhead: Conventional TLS requires symmetric encryption and decryption on each record, which breaks IVSHMEM’s zero-copy path and forces additional data copies and context switches[13][23]. In a high-throughput IVSHMEM environment—where direct page mappings sustain multiple gigabytes per second—this per-record crypto overhead introduces unacceptable latency and CPU load.
  2. 2.
    Limited Shared-Memory Capacity: Unlike network channels, IVSHMEM regions are fixed and small (e.g., 1 MiB)[8][9]. To prevent a malicious VM or service from overwhelming the shared-memory resource, the hypervisor must strictly assign and enforce per-service channel quotas. Conventional socket-based protocols provide no mechanism for hypervisor-driven, size-limited region allocation or fine-grained resource control.
  3. 3.
@@
Conventional end-to-end protocols such as TLS, IPsec, and DTLS are ill-suited for IVSHMEM communication because they (1) impose per-record cryptographic overhead that breaks zero-copy performance[24][23], (2) offer no mechanism to enforce fixed, size-limited shared-memory quotas, and (3) depend on network-layer identifiers (e.g., hostnames, IP addresses) while lacking support for hypervisor-driven channel assignment[14]. Securing IVSHMEM therefore requires a specialized protocol that leverages IVSHMEM’s in-host, fixed-region semantics and hypervisor control, providing end-to-end protection with minimizing additional performance overhead.
IV Design Proposal
@@
IV-A Service-based Channel Separation
Figure 2: Service-based channel separation in Secure IVSHMEM: the trusted host controls metadata in the Control Section and is assigned its own channel in the Data Section for host–VM communication, while each service pair communicates over its dedicated Data Section channel.
@@ IV-A Service-based Channel Separation
An exception to this rule is the first channel in the Data Section. This channel is of a fixed size and is exclusively used for communication between the trusted host and the VMs—for instance, during the initial handshake when a VM sends data to the trusted host. All VMs have access to this channel.
@@
IV-B Granualr Kernel Module Enforcement
@@
IV-B1 Kernel Module Integration
Our kernel module hooks all IVSHMEM-related system calls—including open, read, write, and mmap—as well as any I/O control operations targeting the IVSHMEM device. On each intercepted call, the module retrieves the caller’s service_id and consults the Control Section’s metadata to verify that this service identifier matches the channel being accessed. If the service_id does not correspond to that channel’s assigned service, the module denies the operation. This enforcement ensures that only the authorized host or VM service can interact with its designated shared-memory region.
IV-B2 Channel-Specific Enforcement
@@
IV-C Hypervisor-Mediated Mutual-Authentication Handshake
A secure, hypervisor-mediated handshake is essential for IVSHMEM because it protects against spoofing and impersonation on both sides of the shared-memory channel. In our model, the trusted host (e.g., dom0 in Xen, SOS in ACRN[25], or the host in QEMU/KVM) cannot simply trust that any service presenting a request truly owns its claimed endpoint; similarly, a VM service cannot blindly accept control messages or credentials from the host without risk of impersonation. By embedding a mutual-authentication handshake into the Control Section—that is, having each party present and verify cryptographic credentials tied to its service id—the host first confirms that the requesting service is one of its pre-registered, trusted entities, and then the VM service validates that the host’s response really comes from the genuine hypervisor authority. Only once both directions of identity proof succeed do we allocate a dedicated, zero-copy channel in the Data Section. This two-way validation thwarts malicious actors on either side and ensures end-to-end trust before any IVSHMEM communication occurs.
Our proposed protocol is different from the conventional internet based security protocol in that 1) The trusted host (hypervisor) is not merely a passive participant but is responsible for allocating finite resources and establishing the communication channel and 2) The hypervisor functions as a certification authority (CA)[26], validating service credentials and orchestrating the creation of dedicated secure channels between clients and servers.
@@ IV-C Hypervisor-Mediated Mutual-Authentication Handshake
IV-C1 Protocol Steps (Fig. 3)
Figure 3: Secure IVSHMEM handshake flow: eight steps from Client Hello through Host authorization to establishment of the dedicated shared-memory channel
@@
TABLE I: Key threat scenarios and corresponding Secure IVSHMEM defenses
V Implementation
In this section, we implment Secure IVSHMEM through three components: a kernel module that hooks into the UIO PCI driver[27] to enforce per-channel access control, a user-space OpenSSL-based mutual-authentication handshake over the control page, and a BSD-socket-style library for zero-copy ring-buffer data transfers on authenticated channels.
V-A Kernel-Module Integration via Dynamic Hooks
We build on top of the existing UIO PCI driver for IVSHMEM by inserting a lightweight kernel module that intercepts system calls—mmap(), read() and write()—to enforce per-channel access control.
@@ V-A Kernel-Module Integration via Dynamic Hooks
Using a combination of kprobes and ftrace[28], our module attaches to the IVSHMEM driver’s mmap() entry point. In the hook we:
@@
V-B Handshake Implementation
Figure 4: Credential management: the hypervisor publishes a CA certificate and per-service public keys in the Allowed Service List for Secure IVSHMEM
As illustrated in Figure 4, the mutual-authentication handshake is implemented entirely in user-land using OpenSSL[29] and leverages a IVSHMEM as an in-host transport. At the initial setup, the hypervisor generates a 4096-bit RSA CA key and self-signed certificate. It then builds an allowed-service list that maps each (service_ID, VM_ID) pair to its corresponding public key, and publishes that list together with the CA’s public certificate—read-only—to all VMs via a shared directory. Next, the hypervisor issues 2048-bit RSA key pairs for each registered service, retaining the public key in its allowed-service list and provisioning the private key to the guest VM. Each service in the VM uses this private key to generate its certificate when it performs the handshake. At runtime, host and VM exchange certificates over the IVSHMEM control channel, verify each peer’s certificate against the CA and the allowed-service list, and send signed acknowledgments to confirm mutual credential validation. Once both sides have verified the signature, the VM mark the kernel module that the channel is ESTABLISHED, enabling zero-copy operations under the authenticated session.
V-C Application Library Implementation

### v2

@@
Abstract.
[…]
journalyear: 2025copyright: rightsretainedconference: International Symposium on Memory Systems; October 07–08, 2025; Washington, VA, USAbooktitle: International Symposium on Memory Systems (MemSys ’25), October 07–08, 2025, Washington, VA, USAdoi: 10.1145/3767110.3767119isbn: 979-8-4007-2002-4/25/10
1. Introduction
@@
The automotive industry is rapidly evolving, driven by advances in semiconductor technology that have shifted system architectures from traditional microcontrollers (MCUs) to powerful Systems-on-Chip (SoCs). This evolution not only enhances computational capabilities but also paves the way for Software-Defined Vehicles (SDVs), where flexibility, scalability, and rapid updates are paramount. In SDVs, virtualization technology plays a crucial role by enabling the coexistence of multiple virtual machines (VMs) on a single hardware platform, ensuring isolated yet efficient execution of diverse applications. For example, modern cockpit domain controllers often deploy separate VMs for real-time operations (RTOS) and infotainment systems, which is essential for balancing performance and safety(Jiang et al., 2024; Reinhardt and Morgan, 2014).
Figure 1. A performance comparison between IVSHMEM and other inter-VM communication methods.
@@
Inter-VM communication in these environments is critical. Traditional approaches, such as TCP/UDP over a network stack or even UART-based messaging, often lack sufficient speed and resource efficiency(Ren et al., 2016). Alternative solutions like VirtIO offer a para-virtualized communication mechanism through ring buffers (VirtQueues), but they do not fully leverage the benefits of shared physical memory.(Russell, 2008) IVSHMEM(Inter-VM Shared Memory) addresses these limitations by mapping each VM’s virtualized PCI device to a common physical memory region, allowing rapid data exchange through shared memory(Zhang et al., 2023; Li et al., 2023; Ren et al., 2016). As illustrated in Figure 1, IVSHMEM achieves substantially higher bandwidth and lower latency compared to VirtIO-based networking (Virt-Net) and inter-machine TCP communication over Ethernet, highlighting its superior performance for inter-VM communication. Despite its performance advantages, this method introduces significant security challenges; multiple VMs accessing the same memory space creates vulnerabilities where a compromised or malicious VM could potentially access or modify data belonging to another VM(Sreenivasamurthy and Miller, 2019).
@@
This concern is particularly acute in scenarios where critical systems interact with less secure environments. For instance, when an RTOS communicates with an Android-based infotainment VM, there is a tangible risk that a malicious application within Android might tamper with the shared memory region(Reinhardt and Morgan, 2014). Such tampering could result in attacks ranging from man-in-the-middle to eavesdropping, ultimately compromising system stability and safety.
@@
2. Background
In this section, we examine the IVSHMEM mechanism, outline its challenges, and review recent progress.
2.1. IVSHMEM: Mechanism and Architecture
IVSHMEM is a specialized implementation of shared memory IPC designed for virtualized environments. It emulates a virtual PCI device to expose the shared memory’s base address and size to guest VMs(QEMU Project Developers, [n. d.]; Intel Corporation, [n. d.]). As shown in Figure 2, the hypervisor exposes the shared-memory region to the guest VM, and the UIO kernel driver maps the emulated PCI device, allowing applications to access that memory directly via /dev/uioX. The design leverages the standardized PCI configuration to facilitate memory mapping and efficient communication. Specifically, IVSHMEM utilizes:
@@ 2.1. IVSHMEM: Mechanism and Architecture
The doorbell interrupt mechanism enabled by this configuration allows VMs to notify one another when new data is available, ensuring efficient core utilization and reducing latency in inter-VM communication(Kazmi, 2004).
Figure 2. The overview of IVSHMEM architecture
2.2. Security Concerns of Shared Memory
Shared-memory IPC within a single OS benefits from well-established protections, including file access controls, sandboxing, and security modules such as SELinux or AppArmor. IVSHMEM, however, presents distinct challenges (Spencer et al., 1999; AppArmor, [n. d.]). In a traditional OS environment, the OS enforces strict access controls over shared memory regions, ensuring that only authorized processes can read or write data. However, these protections doesn’t work when multiple, potentially untrusted VMs share the same memory space. A compromised or malicious VM could easily access or tamper with data in the shared region, leading to unauthorized data disclosure, corruption, or even system instability(Sreenivasamurthy and Miller, 2019).
2.3. Secure Communication Over Insecure Channels
The challenge of ensuring secure communication in IVSHMEM environments is analogous to securing communication over the Internet, where multiple parties exchange information over an inherently insecure channel. In network communications, protocols such as TLS rely on key exchange mechanisms, mutual authentication, and end-to-end encryption to safeguard data integrity and confidentiality(Rescorla, 2018; Satapathy et al., 2016). Similarly, secure multi-party communication techniques, such as Diffie-Hellman key exchange and advanced encryption standards, are employed to establish trust even when the channel is compromised(Diffie and Hellman, 2022; Rahman et al., 2014).
In the context of IVSHMEM, the situation is even more complex because multiple services must share the same restricted memory space as a communication channel. This necessitates designing a secure protocol that ensures confidentiality and integrity, similar to TLS or other network security protocols, while also accommodating the shared nature of the memory resource. Our research addresses these challenges by proposing a secure protocol that protect the data transmitted via IVSHMEM, while also mitigating the performance overhead typically associated with such security measures.
2.4. Recent Progress of IVSHMEM Communication
@@
The SIVSHM project introduces a segmentation approach to IVSHMEM, enhancing security by isolating shared memory regions among VMs. However, this strict isolation introduces notable overheads from reduced buffer size. (Sreenivasamurthy and Miller, 2019).
Performance-centric solutions, such as XenLoop and MemPipe, have been proposed to optimize IVSHMEM by improving transparency and reducing latency(Zhang et al., 2007)(Ren et al., 2016)(Zhang and Liu, 2016)(Wang et al., 2008)(Diakhaté et al., 2008)(Macdonell, 2011). These solutions integrate seamlessly with traditional socket-based network stacks, allowing applications to benefit from high-speed shared memory communication without explicit changes. However, these solutions primarily prioritize performance and transparency and lack of mechanisms for security threats(Ren et al., 2016).
Other approaches have leveraged hypervisor-managed policies and features, like Xen’s grant tables, to enforce finer-grained security controls. Grant tables establish explicit, controlled memory-sharing agreements between VMs, restricting access to designated regions. However, this technique is tied to a specific hypervisor and does not provide a general-purpose driver interface (Xen Project, 2025).
@@ 3.4. Security Goals
3.5. Limitations of Conventional Security Protocols
@@
  1. (1)
    Performance Overhead: Conventional TLS requires symmetric encryption and decryption on each record, which breaks IVSHMEM’s zero-copy path and forces additional data copies and context switches(Rescorla, 2018)(Müller et al., 2014). In a high-throughput IVSHMEM environment—where direct page mappings sustain multiple gigabytes per second—this per-record crypto overhead introduces unacceptable latency and CPU load.
  2. (2)
    Limited Shared-Memory Capacity: Unlike network channels, IVSHMEM regions are fixed and small (e.g., 1 MiB)(QEMU Project Developers, [n. d.])(Intel Corporation, [n. d.]). To prevent a malicious VM or service from overwhelming the shared-memory resource, the hypervisor must strictly assign and enforce per-service channel quotas. Conventional socket-based protocols provide no mechanism for hypervisor-driven, size-limited region allocation or fine-grained resource control.
  3. (3)
@@
Conventional end-to-end protocols such as TLS, IPsec, and DTLS are ill-suited for IVSHMEM communication because they (1) impose per-record cryptographic overhead that breaks zero-copy performance(Shue et al., 2005; Müller et al., 2014), (2) offer no mechanism to enforce fixed, size-limited shared-memory quotas, and (3) depend on network-layer identifiers (e.g., hostnames, IP addresses) while lacking support for hypervisor-driven channel assignment(Satapathy et al., 2016). Securing IVSHMEM therefore requires a specialized protocol that leverages IVSHMEM’s in-host, fixed-region semantics and hypervisor control, providing end-to-end protection with minimizing additional performance overhead.
4. Design Proposal
@@
4.1. Service-based Channel Separation
Figure 3. Service-based channel separation in Secure IVSHMEM: the trusted host controls metadata in the Control Section and is assigned its own channel in the Data Section for host–VM communication, while each service pair communicates over its dedicated Data Section channel.
@@ 4.1. Service-based Channel Separation
An exception to this rule is the first channel in the Data Section. This channel is of a fixed-size and is exclusively used for communication between the trusted host and the VMs, such as during the initial handshake when a VM sends data to the trusted host. All VMs have access to this channel.
@@
4.2. Granular Kernel Module Enforcement
@@
4.2.1. Kernel Module Integration
Our kernel module hooks all IVSHMEM-related system calls, including open, read, write, and mmap, as well as any I/O control operations targeting the IVSHMEM device. On each intercepted call, the module retrieves the caller’s service_id and checks the Control Section’s metadata to verify that this service identifier matches the channel being accessed. If the service_id does not correspond to that channel’s assigned service, the module denies the operation. This enforcement ensures that only the authorized host or VM service can interact with its designated shared-memory region.
4.2.2. Channel-Specific Enforcement
@@
4.3. Hypervisor-Mediated Mutual-Authentication Handshake
A secure, hypervisor-mediated handshake is essential for IVSHMEM because it protects against spoofing and impersonation on both sides of the shared-memory channel. In our model, the trusted host (e.g., dom0 in Xen, SOS in ACRN(Li et al., 2019), or the host in QEMU/KVM) cannot simply trust that any service presenting a request truly owns its claimed endpoint; similarly, a VM service cannot blindly accept control messages or credentials from the host without risk of impersonation. By embedding a mutual-authentication handshake into the Control Section—that is, having each party present and verify cryptographic credentials tied to its service id—the host first confirms that the requesting service is one of its pre-registered, trusted entities, and then the VM service validates that the host’s response really comes from the genuine hypervisor authority. Only once both directions of identity proof succeed do we allocate a dedicated, zero-copy channel in the Data Section. This two-way validation thwarts malicious actors on either side and ensures end-to-end trust before any IVSHMEM communication occurs.
Our proposed protocol is different from the conventional internet based security protocol in that 1) The trusted host (hypervisor) is not merely a passive participant but is responsible for allocating finite resources and establishing the communication channel and 2) The hypervisor functions as a certification authority (CA)(Perez et al., 2006), validating service credentials and orchestrating the creation of dedicated secure channels between clients and servers.
@@ 4.3. Hypervisor-Mediated Mutual-Authentication Handshake
4.3.1. Protocol Steps (Fig. 4)
Figure 4. Secure IVSHMEM handshake flow: eight steps from Client Hello through Host authorization to establishment of the dedicated shared-memory channel
@@
Table 1. Key threat scenarios and corresponding Secure IVSHMEM defenses
4.4. Ring Buffer and Anticipation Timing Window
Figure 5. The IVSHMEM protocol is optimized using a Ring Buffer and an Anticipation Time Window mechanism.
To enhance performance in our IVSHMEM protocol design, we adopt two key mechanisms: a ring buffer and an anticipation timing window. Ring buffers are widely used in shared memory communication due to their low-overhead, lock-free design, which supports efficient producer-consumer patterns. This structure allows continuous data exchange without the need for frequent synchronization, thereby improving throughput and reducing contention.
In addition, we incorporate the anticipation timing window mechanism, originally proposed in (Zhang and Liu, 2016), to mitigate the overhead associated with frequent interrupt handling. Rather than triggering an interrupt (IRQ) for every data transmission, which incurs significant system-wide context switching overhead, the receiver VM periodically polls the shared buffer for a predefined time window. This approach reduce interrupt cost while maintaining responsiveness, ultimately reducing communication latency and improving overall efficiency.
The performance benefits of these optimizations are evaluated in detail in Section 6.
5. Implementation
In this section, we implment Secure IVSHMEM through three components: a kernel module that hooks into the UIO PCI driver(Weisbach et al., 2011) to enforce per-channel access control, a user-space OpenSSL-based mutual-authentication handshake over the control page, and a BSD-socket-style library for zero-copy ring-buffer data transfers on authenticated channels.
5.1. Kernel-Module Integration via Dynamic Hooks
We build on top of the existing UIO PCI driver for IVSHMEM by inserting a lightweight kernel module that intercepts system calls(mmap(), read() and write())to enforce per-channel access control.
@@ 5.1. Kernel-Module Integration via Dynamic Hooks
Using a combination of kprobes and ftrace(Gebai and Dagenais, 2018), our module attaches to the IVSHMEM driver’s mmap() entry point. In the hook we:
@@
5.2. Handshake Implementation
Figure 6. Credential management: the hypervisor publishes a CA certificate and per-service public keys in the Allowed Service List for Secure IVSHMEM
As illustrated in Figure 6, the mutual-authentication handshake is implemented entirely in user-land using OpenSSL(Viega et al., 2002) and leverages a IVSHMEM as an in-host transport. During the initial setup, the hypervisor generates a 4096-bit RSA CA key and self-signed certificate. It then builds an allowed-service list that maps each (service_ID, VM_ID) pair to its corresponding public key, and publishes that list together with the CA’s public certificate to all VMs via a shared directory. Next, the hypervisor issues 2048-bit RSA key pairs for each registered service, retaining the public key in its allowed-service list and provisioning the private key to the guest VM. Each service in the VM uses this private key to generate its certificate when it performs the handshake. At runtime, host and VM exchange certificates over the IVSHMEM control channel, verify each peer’s certificate against the CA and the allowed-service list, and send signed acknowledgments to confirm mutual credential validation. Once both sides have verified the signature, the VM mark the kernel module that the channel is ESTABLISHED, enabling zero-copy operations under the authenticated session.
5.3. Application Library Implementation